NOAA SATOPS Morning Report: Friday, November 30 2012


SATOPS Morning Report: Friday, November 30, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 335

GOES
GOES-13 DOY334 15:45z NHEMI, 1601z SHEMI, 16:09Z SHEMI, and 15:46z sounding lost due to an operations error.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N18 DOY3334 10:24z lost the first 40 seconds of GAC 2B and 3B due to a WCDA antenna problem.

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
Due to a NWS network problem GINI east lost all sectors for 1415 UTC and 1430 UTC.

The GINI NPORT West lost all sectors for 1700Z due to a delay in processing.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.